OpenStudy (anonymous):

It says to multiply the number (5x10^3)x(7x10^6) How do i do this without a calculator?

OpenStudy (anonymous):

\[(5*10^3) * (7*10^6)\] Follow your order of operations and solve: \[(5*1000) * (7* 1000000)\] \[(5000)*(7000000)\] \[35000000000\]

OpenStudy (anonymous):

OMG YOUR A LIFE SAVER! its so much easier! Thank you :)

OpenStudy (kainui):

There are no orders of operations here, multiplication works any way you want to go. For instance, (2*3)*5=6*5=30 and 2*(3*5)=2*15=30. So the trick here is just to remember WHAT IS AN EXPONENT? Well, luckily that's not so hard, it just tells you how many times you're multiplying that number together. So we look at the number and write: \[\LARGE (5*10^3)*(7*10^6)=3*5*10^3*10^6\] since 10^3 is really 10*10*10 and 10^6 is just 6 tens multiplied together, multiplying all of those tens together at once is just 9 tens in all right? \[\LARGE 3*5*10^3*10^6=35*10^9\] And since 7*5 =35, we can easily just write it. Then, every time you multiply by 10 you are just adding a zero behind the number, so you can write 35 with 9 zeros behind it, but usually people like scientific notation like this.

OpenStudy (anonymous):

much appreciated :)